# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v2

@@ 1 Introduction
To this end, we propose a new CTDG model named DyGMamba. DyGMamba first leverages a node-level Mamba SSM to encode historical node interactions. Another time-level Mamba SSM is then employed to exploit the edge-specific temporal patterns, where its output is used to dynamically select the critical information from the interaction history. To summarize: (1) We present DyGMamba, the first model using SSMs for CTDG representation learning; (2) DyGMamba demonstrates high efficiency and strong efectiveness in modeling long-term temporal dependencies in CTDGs; (3) Experimental results show that DyGMamba achieves new state-of-the-art on dynamic link prediction over most commonly-used CTDG datasets.
@@ 2.1 Related Work
Dynamic graph representation learning methods can be categorized into two groups, i.e., DTDG and CTDG methods. DTDG methods (Pareja et al., 2020; Goyal et al., 2020; Sankar et al., 2020; You et al., 2022; Li et al., 2024) can only model DTDGs where each of them is represented as a sequence of graph snapshots. Modeling a dynamic graph as graph snapshots requires time discretization and will inevitably cause information loss (Kazemi et al., 2020). To overcome this problem, recent works focus more on developing CTDG methods that treat a dynamic graph as a stream of events, where each event has its own unique timestamp. Some works (Trivedi et al., 2019; Chang et al., 2020) model CTDGs by using temporal point process. Another line of works (Xu et al., 2020; Ma et al., 2020; Wang et al., 2021b; Gravina et al., 2024) designs advanced temporal graph neural networks for CTDGs. Besides, some other methods are developed based on memory networks (Kumar et al., 2019; Rossi et al., 2020; Liu et al., 2022), temporal random walk (Wang et al., 2021c; Jin et al., 2022) and temporal sequence modeling (Cong et al., 2023; Yu et al., 2023; Besta et al., 2023; Tian et al., 2024). Since some real-world CTDGs heavily rely on long-term temporal information for effective learning, a number of works start to study how to build CTDG models that can do long range propagation of information over time (Yu et al., 2023; Gravina et al., 2024). We compare them later in experiments with DyGMamba to show our model outperforms them in capturing long-term temporal information.
@@
Transformer (Vaswani et al., 2017) is a de facto backbone architecture in modern deep learning. However, its self-attention mechanism results in large space and time complexity, making it unsuitable for extremely long sequence modeling (Duman Keles et al., 2023). To address this, many works focus on building structured state space models that scale linearly or near-linearly with input sequence length (Gu et al., 2021; 2022b; 2022a; Smith et al., 2023; Peng et al., 2023; Ma et al., 2023; Gu & Dao, 2023). Most structured SSMs exhibit linear time invariance (LTI), meaning their parameters are not input-dependent and fixed for all time-steps. Gu & Dao (2023) demonstrate that LTI prevents SSMs from effectively selecting relevant information from the input context, which is problematic for tasks requiring context-aware reasoning. To solve this issue, Gu & Dao (2023) proposes S6, also known as Mamba, which uses a selection mechanism to dynamically choose important information from input sequence elements. Selection mechanism involves learning functions that map input data to SSM’s parameters, making Mamba both efficient and effective in modeling language, DNA sequences, and audio. In our work, we show how SSM can be used to achieve effective and efficient representation learning on CTDGs.
@@ 2.2 Preliminaries
A structured SSM, i.e., S4 (Gu et al., 2022b), is inspired by a continuous system described below:
$\mathbf{z}(\tau)^{\prime}=\mathbf{A}\mathbf{z}(\tau)+\mathbf{B}q(\tau),\ \ r(\tau)=\mathbf{C}\mathbf{z}(\tau).$ (1a)
$q(\tau)\in\mathbb{R}$ and $r(\tau)\in\mathbb{R}$ are the 1-dimensional input and output over time $\tau$ [^1] We use $\tau$ rather than $t$ to indicate time in a continuous system to distinguish from the time in CTDGs., respectively. $\mathbf{A}\in\mathbb{R}^{d_{1}\times d_{1}},\mathbf{B}\in\mathbb{R}^{1\times d_{1}},\mathbf{C}\in\mathbb{R}^{d_{1}\times 1}$ are three parameters deciding the system. Based on it, an S4 includes a time-scale parameter $\Delta\in\mathbb{R}$ and discretizes all the parameters to adapt the model to processing sequences recurrently.
$\mathbf{z}_{\tau}=\bar{\mathbf{A}}\mathbf{z}_{\tau-1}+\bar{\mathbf{B}}p_{\tau},\ \ q_{\tau}=\mathbf{C}\mathbf{z}_{\tau},$ (2)
where $\bar{\mathbf{A}}=\text{exp}(\Delta\mathbf{A}),\ \ \bar{\mathbf{B}}=(\Delta\mathbf{A})^{-1}\text{exp}((\Delta\mathbf{A})-\mathbf{I})\Delta\mathbf{B}.$
Here, $\tau$ is also discretized to denote the position of an element in a sequence. When the dimension size of an input $d_{2}$ becomes higher (i.e., $d_{2}>1$), S4 is in a Single-Input Single-Output (SISO) fashion, processing each input dimension in parallel with the same set of parameters. Based on S4, Mamba SSM follows the SISO mode but changes its parameters into input-dependent by employing several trainable linear layers to map input into $\bar{\mathbf{A}}$, $\bar{\mathbf{B}}$ and $\Delta$. The system is evolving as it processes input data sequentially, making Mamba time-variant and suitable for modeling temporal sequences.
@@ 3 DyGMamba
Figure 1 illustrates the overview of DyGMamba. Given a potential interaction $(u,v,t)$, CTDG models are asked to predict whether it exists or not. DyGMamba extracts the historical one-hop interactions of node $u$ and $v$ before timestamp $t$ from the CTDG $\mathcal{G}$ and gets two interaction sequences $\mathcal{S}_{u}^{t}=\{(u,u^{\prime},t^{\prime})|t^{\prime}<t,(u,u^{\prime},t^{\prime})\in\mathcal{G}\}\cup\{(u^{\prime},u,t^{\prime})|t^{\prime}<t,(u^{\prime},u,t^{\prime})\in\mathcal{G}\}$ and $\mathcal{S}_{v}^{t}=\{(v,v^{\prime},t^{\prime})|t^{\prime}<t,(v,v^{\prime},t^{\prime})\in\mathcal{G}\}\cup\{(v^{\prime},v,t^{\prime})|t^{\prime}<t,(v^{\prime},v,t^{\prime})\in\mathcal{G}\}$ containing $u$’s and $v$’s one-hop temporal neighbors $\textit{Nei}_{u}^{t}=\{(u^{\prime},t^{\prime})|(u,u^{\prime},t^{\prime})\text{or }(u^{\prime},u,t^{\prime})\in\mathcal{G},t^{\prime}<t\}$ and $\textit{Nei}_{v}^{t}=\{(v^{\prime},t^{\prime})|(v,v^{\prime},t^{\prime})\text{or }(v^{\prime},v,t^{\prime})\in\mathcal{G}\}$ (link features are omitted for clarity). Then it encodes the neighbors in $\textit{Nei}_{u}^{t}$ and $\textit{Nei}_{v}^{t}$ to get two sequences of encoded neighbor representations for $u$ and $v$. To learn the edge-specific temporal pattern of $(u,v,t)$, we find the interactions between $u$ and $v$ before $t$, compute the time difference between each pair of neighboring interactions, and build a sequence of time differences $\mathcal{S}_{u,v}^{t}$. Finally, DyGMamba dynamically selects critical information from the encoded neighbors and incorporates the learned temporal pattern to achieve link prediction.
@@ 3.1 Learning One-Hop Temporal Neighbors
We first map the padded features of $u$’s and $v$’s one-hop temporal neighbors to the same dimension $d$:
$\mathbf{X}_{\theta}^{t}:=f_{N}(\mathbf{X}_{\theta}^{t}),\ \ \mathbf{E}_{\theta}^{t}:=f_{E}(\mathbf{E}_{\theta}^{t}),\ \ \mathbf{T}_{\theta}^{t}:=f_{T}(\mathbf{T}_{\theta}^{t}),\ \ \mathbf{F}_{\theta}^{t}:=f_{F}(\mathbf{F}_{\theta}^{t}).$ (3)
$f_{N}(\cdot):\mathbb{R}^{p\cdot d_{N}}\to\mathbb{R}^{d}$, $f_{E}(\cdot):\mathbb{R}^{p\cdot d_{E}}\to\mathbb{R}^{d}$, $f_{T}(\cdot):\mathbb{R}^{p\cdot d_{T}}\to\mathbb{R}^{d}$, $f_{F}(\cdot):\mathbb{R}^{p\cdot d_{F}}\to\mathbb{R}^{d}$ are four MLPs for different types of neighbor features. We take the concatenation of them as the encoded representations of the temporal neighbors, i.e., $\mathbf{H}_{\theta}^{t}=\mathbf{X}_{\theta}^{t}\|\mathbf{E}_{\theta}^{t}\|\mathbf{T}_{\theta}^{t}\|\mathbf{F}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d}$. We input $\mathbf{H}_{u}^{t}$ and $\mathbf{H}_{v}^{t}$ separately into a node-level SSM block to learn the temporal dependencies of temporal neighbors. The node-level SSM block consists of $l_{N}$ layers, where each layer is defined as follows (Eq. 4-5). First, we input $\mathbf{H}_{\theta}^{t}$ into a Mamba SSM
@@
$\bar{\mathbf{A}}_{1}=\text{exp}(\Delta_{1}\mathbf{A}_{1}),\ \ \bar{\mathbf{B}}_{1}=(\Delta_{1}\mathbf{A}_{1})^{-1}\text{exp}((\Delta_{1}\mathbf{A}_{1})-\mathbf{I})\Delta_{1}\mathbf{B}_{1};$ (4c)
@@
$\mathbf{W}_{\mathbf{B}_{1}},\mathbf{W}_{\mathbf{C}_{1}}\in\mathbb{R}^{4d\times d_{\text{SSM}}}$ and $\mathbf{W}_{\Delta_{1}}\in\mathbb{R}^{4d\times 1}$. $\bar{\mathbf{A}}_{1},\bar{\mathbf{B}}_{1}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d\times d_{\text{SSM}}}$ are discretized parameters. $\text{Param}_{\Delta_{1}}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d}$ is a parameter defined by Gu & Dao (2023). $\mathbf{I}$ is an identity matrix, and Eq. 4d denotes the SSM forward process similar to Eq. 2. Then we use an MLP $f_{\text{node}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ on SSM’s output
@@ 3.2 Learning from Edge-Specific Temporal Patterns
$\bar{\mathbf{A}}_{2}=\text{exp}(\Delta_{2}\mathbf{A}_{2}),\ \ \bar{\mathbf{B}}_{2}=(\Delta_{2}\mathbf{A}_{2})^{-1}\text{exp}((\Delta_{2}\mathbf{A}_{2})-\mathbf{I})\Delta_{2}\mathbf{B}_{2};$ (6c)
@@
$\mathbf{W}_{\mathbf{B}_{2}},\mathbf{W}_{\mathbf{C}_{2}}\in\mathbb{R}^{\gamma d\times d_{\text{SSM}}}$ and $\mathbf{W}_{\Delta_{2}}\in\mathbb{R}^{\gamma d\times 1}$. $\bar{\mathbf{A}}_{2},\bar{\mathbf{B}}_{2}\in\mathbb{R}^{k\times\gamma d\times d_{\text{SSM}}}$ are discretized parameters. $\text{Param}_{\Delta_{2}}\in\mathbb{R}^{k\times\gamma d}$ is a parameter defined as same as $\text{Param}_{\Delta_{1}}$. In practice, we set $k$ to a number much smaller than $|\textit{Nei}_{\theta}^{t}|$, e.g., $10$. This ensures that time-level SSM will not incur huge computational burden and the model focuses more on the recent histories. Note that we cannot always find $k$ recent interactions between $u$ and $v$. To enable batch processing, we set the time difference without a found historical interaction to a very large number $10^{10}$. For example, if $k=2$, and for $(u,v,t)$ we can only find $(u,v,t_{0})$. The time difference sequence will be $\{10^{10},t-t_{0}\}$. $10^{10}$ is much larger than $t-t_{0}$, indicating that $u$ and $v$ have not had an interaction for an extremely long time, same as existing no historical interaction.
@@
$f_{\text{map}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{4d}$ and $f_{\text{map}3}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{1}$ are two mapping MLPs. $f^{\prime}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ is another MLP introducing training parameters. Note that $\alpha_{u}$/$\alpha_{v}$ is computed by considering both the edge-specific temporal pattern and the opposite node $v$/$u$. In the node-level SSM block, we separately model the one-hop temporal neighbors of each node $\theta$, making it hard to connect $u$ and $v$. Computing $\alpha_{\theta}$ as Eq. 7c helps to strengthen the connection between both nodes and meanwhile incorporates the learned temporal pattern. $\beta_{\theta}$ is derived by transforming the queried results based on $\alpha_{\theta}$ into weights. It is then used to compute a weighted-sum of all temporal neighbors for representing $\theta$ at $t$, i.e., $\mathbf{h}_{\theta}^{t}$. Finally, we output the representations of $u$, $v$ and the edge-specific temporal pattern by employing two output MLPs $f_{\text{out}1}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{d_{N}}$ and $f_{\text{out}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{d_{N}}$
$\mathbf{h}_{\theta}^{t}:=f_{\text{out}1}(\mathbf{h}_{\theta}^{t})\in\mathbb{R}^{d_{N}},\ \mathbf{h}_{u,v}^{t}:=f_{\text{out}2}(\mathbf{h}_{u,v}^{t})\in\mathbb{R}^{d_{N}}.$ (8)

### v3

@@ 1 Introduction
To this end, we propose a new CTDG model named DyGMamba. DyGMamba first leverages a node-level Mamba SSM to encode historical node interactions. Another time-level Mamba SSM is then employed to exploit the edge-specific temporal patterns, where its output is used to dynamically select the critical information from the interaction history. To summarize: (1) We present DyGMamba, the first model using SSMs for CTDG representation learning; (2) DyGMamba demonstrates high efficiency and strong effectiveness in modeling long-term temporal dependencies in CTDGs; (3) Experimental results show that DyGMamba achieves new state-of-the-art on dynamic link prediction over most commonly-used CTDG datasets.
@@ 2.1 Related Work
Dynamic graph representation learning methods can be categorized into two groups, i.e., DTDG and CTDG methods. DTDG methods (Pareja et al., 2020; Goyal et al., 2020; Sankar et al., 2020; You et al., 2022; Li et al., 2024a) can only model DTDGs where each of them is represented as a sequence of graph snapshots. Modeling a dynamic graph as graph snapshots requires time discretization and will inevitably cause information loss (Kazemi et al., 2020). To overcome this problem, recent works focus more on developing CTDG methods that treat a dynamic graph as a stream of events, where each event has its own unique timestamp. Some works (Trivedi et al., 2019; Chang et al., 2020) model CTDGs by using temporal point process. Another line of works (Xu et al., 2020; Ma et al., 2020; Wang et al., 2021b; Gravina et al., 2024) designs advanced temporal graph neural networks for CTDGs. Besides, some other methods are developed based on memory networks (Rossi et al., 2020; Liu et al., 2022), temporal random walk (Wang et al., 2021c; Jin et al., 2022) and temporal sequence modeling (Cong et al., 2023; Yu et al., 2023; Tian et al., 2024). Since some real-world CTDGs heavily rely on long-term temporal information for effective learning, a number of works start to develop CTDG models that can do long range propagation of information over time (Yu et al., 2023; Gravina et al., 2024).
@@
Transformer (Vaswani et al., 2017) is a de facto backbone architecture in modern deep learning. However, its self-attention mechanism results in large space and time complexity, making it unsuitable for extremely long sequence modeling (Duman Keles et al., 2023). To address this, many works focus on building structured state space models that scale linearly or near-linearly with input sequence length (Gu et al., 2021; 2022b; 2022a; Smith et al., 2023; Peng et al., 2023; Ma et al., 2023; Gu & Dao, 2023). Most structured SSMs exhibit linear time invariance (LTI), meaning their parameters are not input-dependent and fixed for all time-steps. Gu & Dao (2023) demonstrate that LTI prevents SSMs from effectively selecting relevant information from the input context, which is problematic for tasks requiring context-aware reasoning. To solve this issue, Gu & Dao (2023) proposes S6, also known as Mamba, which uses a selection mechanism to dynamically choose important information from input sequence elements. Selection mechanism involves learning functions that map input data to SSM’s parameters, making Mamba both efficient and effective in modeling language, DNA sequences, and audio. Recently, there have been several works employing Mamba SSM for representation learning on static graphs (Wang et al., 2024; Behrouz & Hashemi, 2024) and spatial-temporal graphs (Li et al., 2024b). Unlike our work, they do not focus on capturing the dynamic evolution of graph structure and are thus not suitable for CTDG modeling.
@@ 2.2 Preliminaries
S4 and Mamba (Gu et al., 2022b; Gu & Dao, 2023) are inspired by a continuous system which can be described as $\mathbf{z}(\tau)^{\prime}=\mathbf{A}\mathbf{z}(\tau)+\mathbf{B}q(\tau)$ and $r(\tau)=\mathbf{C}\mathbf{z}(\tau)$. $q(\tau)\in\mathbb{R}$ and $r(\tau)\in\mathbb{R}$ are the 1-dimensional input and output over time $\tau$ [^1] We use $\tau$ rather than $t$ to indicate time in a continuous system to distinguish from the time in CTDGs., respectively. $\mathbf{A}\in\mathbb{R}^{d_{1}\times d_{1}},\mathbf{B}\in\mathbb{R}^{d_{1}\times 1},\mathbf{C}\in\mathbb{R}^{1\times d_{1}}$ are three parameters deciding the system. Based on it, both S4 and Mamba include a time-scale parameter $\Delta\in\mathbb{R}$ and discretize all the parameters to adapt to a discretized system
$\mathbf{z}_{\tau}=\bar{\mathbf{A}}\mathbf{z}_{\tau-1}+\bar{\mathbf{B}}p_{\tau},\ \ q_{\tau}=\mathbf{C}\mathbf{z}_{\tau};\ \bar{\mathbf{A}}=\text{exp}(\Delta\mathbf{A}),\ \bar{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\text{exp}(\Delta\mathbf{A})-\mathbf{I})\Delta\mathbf{B}.$ (1)
Here, $\tau$ is also discretized to denote the position of a sequence element. Given Eq. 1 , sequence processing with S4 and Mamba can be written as computing an output sequence with convolution
$\mathbf{q}=\mathbf{p}*\bar{\mathbf{K}}_{\text{SSM}},\ \text{where}\ \bar{\mathbf{K}}_{\text{SSM}}=[\mathbf{C}\bar{\mathbf{B}},\mathbf{C}\bar{\mathbf{A}}\bar{\mathbf{B}},...,\mathbf{C}\bar{\mathbf{A}}^{|\mathbf{p}|-1}\bar{\mathbf{B}}]\in\mathbb{R}^{|\mathbf{p}|-1}.$ (2)
$\mathbf{p}\in\mathbb{R}^{|\mathbf{p}|}$ and $\mathbf{q}\in\mathbb{R}^{|\mathbf{p}|}$ are input and output sequences, where $|\mathbf{p}|$ is the sequence length of $\mathbf{p}$. * denotes the element-wise multiplication. When the dimension size of each element $p_{\tau}$ in $\mathbf{p}$ becomes higher (i.e., $p_{\tau}\in\mathbb{R}^{d_{2}}$ is a vector and $d_{2}>1$), both S4 and Mamba are in a Single-Input Single-Output (SISO) fashion, processing each input dimension in parallel with the same set of parameters. We follow Gu & Dao (2023) and denote the computation in Eq. 2 on the input sequences with vector elements as a function $\text{SSM}_{\bar{\mathbf{A}},\bar{\mathbf{B}},\mathbf{C}}(\cdot)$222 Input and output of $\text{SSM}_{\bar{\mathbf{A}},\bar{\mathbf{B}},\mathbf{C}}(\cdot)$ are matrices where each row is a vector corresponding to an element. See App. H for more details of SISO and the function.. Different from S4 which uses same parameters to process each element, Mamba changes its parameters into input-dependent by employing several trainable linear layers to map input into $\bar{\mathbf{B}}$, $\mathbf{C}$ and $\Delta$. The system is evolving as it processes different elements in the input sequence, making Mamba time-variant and suitable for modeling temporal sequences.
@@ 3 DyGMamba
Fig. 1 illustrates the overview of DyGMamba. Given a potential interaction $(u,v,t)$, CTDG models are asked to predict whether it exists or not. DyGMamba extracts the historical one-hop interactions of node $u$ and $v$ before timestamp $t$ from the CTDG $\mathcal{G}$ and gets two interaction sequences $\mathcal{S}_{u}^{t}=\{(u,u^{\prime},t^{\prime})|t^{\prime}<t,(u,u^{\prime},t^{\prime})\in\mathcal{G}\}\cup\{(u^{\prime},u,t^{\prime})|t^{\prime}<t,(u^{\prime},u,t^{\prime})\in\mathcal{G}\}$ and $\mathcal{S}_{v}^{t}=\{(v,v^{\prime},t^{\prime})|t^{\prime}<t,(v,v^{\prime},t^{\prime})\in\mathcal{G}\}\cup\{(v^{\prime},v,t^{\prime})|t^{\prime}<t,(v^{\prime},v,t^{\prime})\in\mathcal{G}\}$ containing $u$’s and $v$’s one-hop temporal neighbors $\textit{Nei}_{u}^{t}=\{(u^{\prime},t^{\prime})|(u,u^{\prime},t^{\prime})\text{or }(u^{\prime},u,t^{\prime})\in\mathcal{G},t^{\prime}<t\}$ and $\textit{Nei}_{v}^{t}=\{(v^{\prime},t^{\prime})|(v,v^{\prime},t^{\prime})\text{or }(v^{\prime},v,t^{\prime})\in\mathcal{G}\}$ (link features are omitted for clarity). Then it encodes the neighbors in $\textit{Nei}_{u}^{t}$ and $\textit{Nei}_{v}^{t}$ to get two sequences of encoded neighbor representations for $u$ and $v$. To learn the edge-specific temporal pattern of $(u,v,t)$, we find the interactions between $u$ and $v$ before $t$, compute the time difference between each pair of neighboring interactions, and build a sequence of time differences $\mathcal{S}_{u,v}^{t}$. Finally, DyGMamba dynamically selects critical information by assigning different weights to different encoded neighbors based on the learned temporal pattern, and uses the selected information to achieve link prediction.
@@ 3.1 Learning One-Hop Temporal Neighbors
We first map the padded features of $u$’s and $v$’s one-hop temporal neighbors to the same dimension $d$, i.e., $\mathbf{X}_{\theta}^{t}:=f_{N}(\mathbf{X}_{\theta}^{t})$, $\mathbf{E}_{\theta}^{t}:=f_{E}(\mathbf{E}_{\theta}^{t})$, $\mathbf{T}_{\theta}^{t}:=f_{T}(\mathbf{T}_{\theta}^{t})$, $\mathbf{F}_{\theta}^{t}:=f_{F}(\mathbf{F}_{\theta}^{t})$. $f_{N}(\cdot):\mathbb{R}^{p\cdot d_{N}}\to\mathbb{R}^{d}$, $f_{E}(\cdot):\mathbb{R}^{p\cdot d_{E}}\to\mathbb{R}^{d}$, $f_{T}(\cdot):\mathbb{R}^{p\cdot d_{T}}\to\mathbb{R}^{d}$, $f_{F}(\cdot):\mathbb{R}^{p\cdot d_{F}}\to\mathbb{R}^{d}$ are four MLPs for different types of neighbor features. We take the concatenation of them as the encoded representations of the temporal neighbors, i.e., $\mathbf{H}_{\theta}^{t}=\mathbf{X}_{\theta}^{t}\|\mathbf{E}_{\theta}^{t}\|\mathbf{T}_{\theta}^{t}\|\mathbf{F}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d}$. We input $\mathbf{H}_{u}^{t}$ and $\mathbf{H}_{v}^{t}$ separately into a node-level SSM block to learn the temporal dependencies of temporal neighbors. The node-level SSM block consists of $l_{N}$ layers, where each layer is defined as follows (Eq. 3-4). First, we input $\mathbf{H}_{\theta}^{t}$ into a Mamba SSM
@@
$\bar{\mathbf{A}}_{1}=\text{exp}(\Delta_{1}\mathbf{A}_{1}),\ \ \bar{\mathbf{B}}_{1}=(\Delta_{1}\mathbf{A}_{1})^{-1}(\text{exp}(\Delta_{1}\mathbf{A}_{1})-\mathbf{I})\Delta_{1}\mathbf{B}_{1};$ (3c)
@@
$\mathbf{W}_{\mathbf{B}_{1}},\mathbf{W}_{\mathbf{C}_{1}}\in\mathbb{R}^{4d\times d_{\text{SSM}}}$ and $\mathbf{W}_{\Delta_{1}}\in\mathbb{R}^{4d\times 1}$. $\bar{\mathbf{A}}_{1},\bar{\mathbf{B}}_{1}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d\times d_{\text{SSM}}}$ are discretized parameters. $\text{Param}_{\Delta_{1}}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d}$ is a parameter defined by Gu & Dao (2023). $\text{Broadcast}_{4d}(\cdot)$ is a function that copies its vector input for $4d$ times to form a matrix with $4d$ identical columns (following the definition in Gu & Dao (2023) ). $\mathbf{I}$ is an identity matrix. Then we use an MLP $f_{\text{node}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ on SSM’s output
@@ 3.2 Learning from Edge-Specific Temporal Patterns
$\bar{\mathbf{A}}_{2}=\text{exp}(\Delta_{2}\mathbf{A}_{2}),\ \ \bar{\mathbf{B}}_{2}=(\Delta_{2}\mathbf{A}_{2})^{-1}(\text{exp}(\Delta_{2}\mathbf{A}_{2})-\mathbf{I})\Delta_{2}\mathbf{B}_{2};$ (5c)
@@
$\mathbf{W}_{\mathbf{B}_{2}},\mathbf{W}_{\mathbf{C}_{2}}\in\mathbb{R}^{\gamma d\times d_{\text{SSM}}}$ and $\mathbf{W}_{\Delta_{2}}\in\mathbb{R}^{\gamma d\times 1}$. $\bar{\mathbf{A}}_{2},\bar{\mathbf{B}}_{2}\in\mathbb{R}^{k\times\gamma d\times d_{\text{SSM}}}$ are discretized parameters. $\text{Param}_{\Delta_{2}}\in\mathbb{R}^{k\times\gamma d}$ is a parameter defined as same as $\text{Param}_{\Delta_{1}}$. In practice, we set $k$ to a number much smaller than $|\textit{Nei}_{\theta}^{t}|$, e.g., $10$. This ensures that time-level SSM will not incur huge computational burden and the model focuses more on the recent histories. Note that we cannot always find $k$ recent historical interactions between each pair of nodes, leading to varying lengths of time difference sequences for different $(u,v,t)$ in a batch of data. To enable batch processing, we set the time difference without a found historical interaction to a very large number $10^{10}$. For example, if $k=2$, and for $(u,v,t)$ we can only find $(u,v,t_{0})$. The time difference sequence will be $\{10^{10},t-t_{0}\}$. $10^{10}$ is much larger than $t-t_{0}$, indicating that $u$ and $v$ have not had an interaction for an extremely long time, same as existing no historical interaction. We further explain why we use SSM to learn temporal patterns in App. I.
@@
$f_{\text{map}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{4d}$ and $f_{\text{map}3}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{1}$ are two mapping MLPs. $f^{\prime}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ is another MLP introducing training parameters. Note that $\alpha_{u}$/$\alpha_{v}$ is computed by considering both the edge-specific temporal pattern and the opposite node $v$/$u$. In the node-level SSM block, we separately model the one-hop temporal neighbors of each node $\theta$, making it hard to connect $u$ and $v$. Computing $\alpha_{\theta}$ as Eq. 6c helps to strengthen the connection between both nodes and meanwhile incorporates the learned temporal pattern. $\beta_{\theta}$ is derived by transforming the queried results based on $\alpha_{\theta}$ into weights. It is then used to compute a weighted-sum of all temporal neighbors for representing $\theta$ at $t$, i.e., $\mathbf{h}_{\theta}^{t}$. The neighbors assigned with greater weights from $\beta_{\theta}$ are selected as more critical and will contribute more to $\mathbf{h}_{\theta}^{t}$. Finally, we output the representations of $u$, $v$ and the edge-specific temporal pattern by employing two output MLPs $f_{\text{out}1}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{d_{N}}$ and $f_{\text{out}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{d_{N}}$, i.e., $\mathbf{h}_{\theta}^{t}:=f_{\text{out}1}(\mathbf{h}_{\theta}^{t})\in\mathbb{R}^{d_{N}}$, $\mathbf{h}_{u,v}^{t}:=f_{\text{out}2}(\mathbf{h}_{u,v}^{t})\in\mathbb{R}^{d_{N}}$.